OpenStudy (anonymous):

Describe the relationship between the weight of water in the catch bucket and the object in the beaker. A. The weight of the volume of the water is less than the buoyant force of the object. B. The weight of the volume of the water is equal to the buoyant force of the object. C. The weight of the volume of the water is greater than the buoyant force of the object. D. The weight of the volume of the water is not a factor in the buoyant force of the object. @jerr1996 @SamuelAlden917

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

I think it's a

OpenStudy (anonymous):

can u help me another one plzz ;l)

OpenStudy (anonymous):

Yes. And it is A.

OpenStudy (anonymous):

Which of the following will sink in a container of mercury with a density of 13.6 g/mL? A. zinc, D = 7.14 g/mL B. ice, D = 0.92 g/mL C. gold, D = 19.3 g/mL D. gasoline, D = 0.69 g/mL

OpenStudy (anonymous):

c

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

Let me see.

OpenStudy (anonymous):

u answered that pretty fast

OpenStudy (anonymous):

ok